Output is produced according to Q=4LK, where L is the quantity of labor input and K is the quantity of capital input. If the price of K is $10 and the price of L is $5,then the cost minimizing combination of K and L capable of producing 32 units of output is:
Answer #1

Cost = 5*L+10*K

Q=4LK

32 = 4LK


L= 8/K


Cost = 5*L+10*K= 5*8/K +10K

For cost minimization

dCost/dK = -40/K^2 +10

K= 2

L= 8/2= 4

K=2

L=4